# Network Oblivious Transfer via Noisy Channels: Limits and Capacities

### v2

@@ Abstract
In this paper, we aim to study the information-theoretical limits of oblivious transfer. This work also investigates the problem of oblivious transfer over a noisy multiple access channel involving two non-colluding senders and a single receiver. The channel model is characterized by correlations among the parties, with the parties assumed to be either honest-but-curious or, in the receiver’s case, potentially malicious. At first, we study the information-theoretical limits of oblivious transfer between two parties and extend it to the multiple access channel model. We propose a multiparty protocol for honest-but-curious parties where the general multiple access channel is reduced to a certain correlation. In scenarios where the receiver is malicious, the protocol achieves an achievable rate region.
@@ I Introduction
Oblivious Transfer (OT), a fundamental primitive in secure multiparty computation (MPC), plays a central role in the design of cryptographic protocols. OT is complete in the sense that, given access to an OT protocol between two parties, any function can be securely computed between them. More precisely, if OT is assumed to be a trusted primitive available between two parties, then it suffices to achieve general secure computation, even in the presence of malicious adversaries [1]. The most basic form of OT is the 11-out-of-22 OT, where a sender (Alice) holds two distinct messages and a receiver (Bob) selects one of them to receive, without revealing his choice to Alice. At the same time, Bob learns nothing about the unchosen message. A natural extension is the 11-out-of-$m$ OT, in which Alice holds $m$ messages (e.g., bit strings), and Bob retrieves exactly one, without gaining information about the others or revealing his selection to Alice.
@@
OT was first introduced by Rabin [2]. In Rabin’s form, Alice sends a message to Bob with the probability of $\frac{1}{2}$ while she remains oblivious to whether or not Bob received the message. This model is called the ”Erasure Channel” with the erasure probability equal $\frac{1}{2}$. After that, a basic OT protocol was introduced by Even, Goldreich, and Lempel (EGL) [3]. It is well-known that achieving multi-party security (as a basic model) in noise-free communication is impossible [4]. It has been shown that achieving two-party secure communication over a noiseless channel is possible by randomness sharing [5, 6]. Shared randomness consists of random variables known to all communicating parties but independent of the message being transmitted [7].
@@
Up to now, some primary channels have been studied for OT purposes obtained from noise. The binary symmetric channel (BSC) has made a more outstanding contribution [8, 9, 10]. However, all the cryptogates and channels that can be used for obtaining OT are characterized by Kilian in the case of passive adversaries [11]. It should be mentioned that most of the research works in this field have been done from the perspective of the basics of cryptography, and there are a few sources that study the problem of OT from the information theory point of view. However, the OT capacity of noisy channels is generally unknown. It is known to be non-negligible if the players (sometimes we call senders and receivers as players/parties) are committed to the protocol and implement it faithfully, not turning away from additional information (honest-but-curious players). Still, in the case of fully malicious players (active adversaries), non-zero rates have not ever been achieved [12].
@@
As the first step, Nascimento and Winter study the OT capacity of noisy correlations [13], wherein they characterized which noisy channels and distributions are useful for obtaining OT. In [6], they showed that for honest-but-curious players, the OT capacity of noisy resources is positive by achieving a lower bound that coincides with the upper bound of [5]. The OT capacity of the binary erasure channel (BEC) is studied in [14], in which the OT capacity is $C_{\text{OT}}=\frac{1}{2}$ with erasure probability $\frac{1}{2}$ that is a property of the channel/system model in the case of honest-but-curious players and a lower bound is calculated in the case of fully malicious players. Ahlswede and Csiszar achieved a lower and upper bound on the OT capacity of noisy channels [5]. The upper bound is general and valid for every noisy channel with honest-but-curious players, while the lower bound is just valid for a special reduced version of a DMC, wherein the channel outputs are separable into two distinct sets: fully erased bits and fully received bits. It can be easily seen that the upper bound in [5] and the lower bound in [6] for a special reduced version coincide. An improved upper bound compared to [5] is proved in [15] based on a monotonicity property of the tension region in the channel model.
@@
The limits of OT are another important bottleneck in secure two-party computation, so that is addressed in various cryptographic-based papers [16, 17]. Some limitations relate to the capability of cheating parties, and some of them relate to the power of noise over the communication channels [18]. The information-theoretical limits of OT between two parties over a shared channel between three nodes are first addressed in [19] from the same authors. In this paper, we have a more general contribution. As a motivation, we would like to draw attention to the following statement by Kurt Gödel [20]:
@@ 4. Information Theoretic Definitions
Let $P_{UVW}$ be a probability distribution over $\mathcal{U}\times\mathcal{V}\times\mathcal{W}$ For any $\epsilon>0$ and $\epsilon^{\prime}>0$ it holds that [21]:
@@
Also, $H^{\epsilon}_{\infty}(U,V|W)$ can be bounded from below and above as [21]:
@@
[21, Corollary 2.12] Let $P_{X^{n}Y^{n}}$ be independent and identically distributed (i.i.d.) according to $P_{XY}$ over the alphabet $\mathcal{X}^{n}\times\mathcal{Y}^{n}$. For any $\epsilon>0$, we have
@@
A function $h:\mathcal{R}\times\mathcal{X}\to\{0,1\}^{n}$ is a two-universal hash function [22] if, for any $x_{0}\neq x_{1}\in\mathcal{X}$ and for $R$ uniformly distributed over $\mathcal{R}$, it holds that
@@
[6, 23] (Distributed leftover hash lemma) Let $\epsilon>0$, $\epsilon^{\prime}\geq 0$, and let $g_{i}:\mathcal{T}_{i}\times\mathcal{X}_{i}\to\{0,1\}^{n_{i}}$ for $1\leq i\leq m$ be two-universal hash functions. Assume random variables $X_{i}$ over $\mathcal{X}_{i}$, $1\leq i\leq m$, where for any subset $\mathcal{S}\subseteq\{1,2,\dots,m\}$, and $\mathbf{X}|_{\mathcal{S}}=X_{\mathcal{S}(1)},X_{\mathcal{S}(2)},\dots,X_{\mathcal{S}(|\mathcal{S}|)}$, we have
@@
We also briefly note that following from [24, Th. 17.3.3], it directly follows from the previously defined $X_{i}$, $T_{i}$, $g_{i}(T_{i},X_{i})$, and $Z$ that
@@
[25, Corollary 4] Let $p_{X,Y}$ be any probability distribution, where $X\in\mathcal{X}$, $Y\in\mathcal{Y}$, and $y$ is a specific realization of $Y$. Assume that $H_{2}(X|Y=y)\geq c$ for some constant $c$. Let $\mathcal{K}$ be a universal class of functions mapping $X$ to $\{0,1\}^{l}$, and let $\kappa$ be sampled uniformly from $\mathcal{K}$. Then:
@@
[5, Lemma 3] Let $X$, $Y$, and $Z$ be random variables defined on the finite sets $\mathcal{X}$, $\mathcal{Y}$, and $\mathcal{Z}$, respectively. For any $z_{1},z_{2}\in\mathcal{Z}$ with $p\triangleq\Pr[Z=z_{1}]>0$ and $q\triangleq\Pr[Z=z_{2}]>0$, the following inequality holds:
@@ III-A1 Yao’s Millionaires’ Problem
One of the foundational problems that introduced the concept of secure two-party computation is Yao’s Millionaires’ Problem. In this thought experiment, proposed by Andrew Yao in 1982, two millionaires wish to determine who is richer without revealing their actual wealth to each other. Formally, each party holds a private input—Alice has $x$ and Bob has yy—and they aim to compute the function $f(x,y)=(x>y)$ without disclosing any other information about $x$ or $y$. To solve this problem, Yao introduced the garbled circuit technique, where one party constructs a Boolean circuit representing the function and encrypts (or “garbles”) its components. The other party evaluates the garbled circuit using encrypted inputs obtained via oblivious transfer, ultimately learning only the output. This protocol demonstrated the feasibility of general secure computation between two parties and laid the groundwork for modern cryptographic protocols in this domain [26].
@@ III-A2 Blum’s Fair Coin Tossing
Blum’s seminal fair coin tossing (FCT) protocol [17] represents one of the earliest cryptographic solutions for enabling two distrustful parties to agree on a random bit in a fair manner. Consider a scenario where Alice and Bob must decide who gets to use their shared home office during a scheduling conflict. Since neither party trusts the other to flip a coin honestly, they employ a cryptographic mechanism to simulate a fair coin toss. Each party independently selects a random value and commits to it using a secure hash function, effectively locking in their choice without revealing it. Once both commitments are exchanged (e.g., over a public channel), they disclose their original values. The XOR or sum of the two values determines the winner: for example, if the result is even, Alice wins; if odd, Bob does. Crucially, the commitment phase prevents either party from altering their input based on the other’s choice, thereby ensuring fairness. If one party refuses to open their commitment or submits an invalid reveal, they automatically lose the toss, which discourages dishonest behavior. This mechanism exemplifies rational fairness in distributed systems, where the optimal strategy for both participants is to follow the protocol faithfully.
Although this game-theoretic interpretation of fairness aligns naturally with rational behavior, it has received comparatively little attention in the mainstream literature on secure MPC [26, 27, 28]. Instead, the field has gravitated toward a more rigorous and adversarially robust notion known as unbiasability. Under this definition, the protocol must guarantee that no coalition of malicious participants can skew the outcome of the computation (in particular, a coin toss) regardless of their strategy. Blum’s original coin-tossing scheme does not meet this stronger requirement; while it prevents a player from biasing the outcome in their favor, it does not fully eliminate the possibility of adversarial influence. The concept of unbiasability has been extensively investigated in cryptographic research, where it is well-established that achieving it requires an honest majority among participants [28, 29]. Notably, Cleve’s impossibility result [16] proves that in the presence of a dishonest minority comprising half or more of the parties, unbiasable coin tossing becomes fundamentally unachievable.
@@ III-A3 Cleve’s Impossibility Result for Fair Coin Tossing
As Cleve’s analysis demonstrates, a malicious Alice has limited capacity to influence the outcome, since she lacks access to Bob’s random bits $x_{i}$ during the protocol execution. Due to the one-way nature of the trapdoor function $F$, and her inability to invert $F_{K_{i}}(x_{i})$ without the corresponding trapdoor $T_{i}$, she cannot predict or bias the XOR outcomes before the trapdoors are revealed. On the other hand, a malicious Bob has more leverage: he may abort the protocol prematurely by withholding the private key $T_{i}$ at some round $i$. In response, Alice replaces the unknown values $x_{i},x_{i+1},\ldots,x_{r}$ with random bits, thereby introducing uncertainty into the final majority computation. However, Cleve shows that even in the worst-case scenario, this form of selective aborting can only shift the outcome of the majority vote by a small amount—specifically, the bias introduced is upper bounded by $O\left(\frac{1}{\sqrt{r}}\right)$. Moreover, he provides a lower bound of $\frac{1}{2r}$, establishing that some nonzero bias is unavoidable in any finite-round protocol. This result implies that perfect fairness—defined as zero bias—is unattainable for two-party coin-tossing protocols with a finite number of rounds. As the number of rounds $r$ increases, the bias diminishes but never vanishes entirely.
@@ III-B1 In praise of noise
Claude Shannon’s Noisy Channel Coding Theorem [30] demonstrates that reliable communication is achievable even in the presence of noise, provided the information transmission rate remains below a critical limit known as channel capacity. This foundational result reveals that noise, while disruptive, is not entirely detrimental; rather, it can be mitigated through appropriate encoding schemes and the strategic use of redundancy. Shannon’s insight challenges the notion that randomness and disorder are purely negative phenomena, illustrating that they can be harnessed within structured frameworks to preserve the integrity of information.
@@
In a broader conceptual sense, entropy serves as a metaphor for disorder or chaos in the universe, with far-reaching implications for how we understand complexity and the emergence of order. Information theory’s treatment of noise thus intersects with deeper philosophical considerations about the relationship between chaos and structure. Michel Serres, in particular, explores this theme through the metaphor of the parasite [31], suggesting that noise is not merely disruptive but can serve as a generative force. According to Serres, noise introduces feedback and transformation within communicative systems, acting as a catalyst for innovation and the emergence of new meanings. Importantly, one of the most profound insights drawn from noise is its role in enabling the possibility of secure communication. In the sense of OT, now we know why OT can not be obtained from scratch and the existence of noisy resources is a crucial condition for OT, as well as other cryptographic primitives. When no noisy resource is available, the system lacks the necessary randomness to obfuscate the client’s actions from the servers. This absence of noise leads to a situation of complete determinism, where all parties can eventually deduce the hidden information, violating the principles of OT.

### v4

@@ Abstract
In this paper, we study the information-theoretic limits of oblivious transfer via noisy channels. We also investigate oblivious transfer over a noisy multiple-access channel with two non-colluding senders and a single receiver. The channel is modeled through correlations among the parties, who may be honest-but-curious or, in the case of the receiver, potentially malicious. We first revisit the information-theoretic limits of two-party oblivious transfer and then extend these results to the multiple-access setting. For honest-but-curious participants, we introduce a multiparty protocol that reduces a general multiple access channel to a suitable correlation model. In scenarios with a malicious receiver, we characterize an achievable oblivious transfer rate region.
@@ I Introduction
Oblivious Transfer (OT), a fundamental primitive in secure multiparty computation (MPC), plays a central role in the design of cryptographic protocols. OT is complete in the sense that, given access to an OT protocol between two parties, any function can be securely computed between them. More precisely, if OT is assumed to be a trusted primitive available between two parties, then it suffices to achieve general secure computation, even in the presence of malicious adversaries [Gold]. The most basic form of OT is the 11-out-of-22 OT, where a sender (Alice) holds two distinct messages and a receiver (Bob) selects one of them to receive, without revealing his choice to Alice. At the same time, Bob learns nothing about the unchosen message. A natural extension is the 11-out-of-$m$ OT, in which Alice holds $m$ messages (e.g., bit strings), and Bob retrieves exactly one, without gaining information about the others or revealing his selection to Alice.
@@
OT was first introduced by Rabin [Rabin1]. In Rabin’s form, Alice sends a message to Bob with the probability of $\frac{1}{2}$ while she remains oblivious to whether or not Bob received the message. This model is called the ”Erasure Channel” with the erasure probability equal $\frac{1}{2}$. After that, a basic OT protocol was introduced by Even, Goldreich, and Lempel (EGL) [EGL]. It is well-known that achieving multi-party security (as a basic model) in noise-free communication is impossible [Kilian1]. It has been shown that achieving two-party secure communication over a noiseless channel is possible by randomness sharing [Rudolf1, Winter1]. Shared randomness consists of random variables known to all communicating parties but independent of the message being transmitted [Rudolf2].
@@
Up to now, some primary channels have been studied for OT purposes obtained from noise. The binary symmetric channel (BSC) has made a more outstanding contribution [Crepeau1, Crepeau2, Stebila]. However, all the cryptogates and channels that can be used for obtaining OT are characterized by Kilian in the case of passive adversaries [Kilian2]. It should be mentioned that most of the research works in this field have been done from the perspective of the basics of cryptography, and there are a few sources that study the problem of OT from the information theory point of view. However, the OT capacity of noisy channels is generally unknown. It is known to be non-negligible if the players (sometimes we call senders and receivers as players/parties) are committed to the protocol and implement it faithfully, not turning away from additional information (honest-but-curious players). Still, in the case of fully malicious players (active adversaries), non-zero rates have not ever been achieved [Crepeau3].
@@
As the first step, Nascimento and Winter study the OT capacity of noisy correlations [Winter3], wherein they characterized which noisy channels and distributions are useful for obtaining OT. In [Winter1], they showed that for honest-but-curious players, the OT capacity of noisy resources is positive by achieving a lower bound that coincides with the upper bound of [Rudolf1]. The OT capacity of the binary erasure channel (BEC) is studied in [Imai], in which the OT capacity is $C_{\text{OT}}=\frac{1}{2}$ with erasure probability $\frac{1}{2}$ that is a property of the channel/system model in the case of honest-but-curious players and a lower bound is calculated in the case of fully malicious players. Ahlswede and Csiszar achieved a lower and upper bound on the OT capacity of noisy channels [Rudolf1]. The upper bound is general and valid for every noisy channel with honest-but-curious players, while the lower bound is just valid for a special reduced version of a DMC, wherein the channel outputs are separable into two distinct sets: fully erased bits and fully received bits. It can be easily seen that the upper bound in [Rudolf1] and the lower bound in [Winter1] for a special reduced version coincide. An improved upper bound compared to [Rudolf1] is proved in [Rao] based on a monotonicity property of the tension region in the channel model. Also, symmetric private information retrieval (SPIR) has been studied under the terminology OT, with a noisy channel between the server and the client to achieve information-theoretic security, e.g., [amir1, amir2, amir3].
@@
The limits of OT are another important bottleneck in secure two-party computation, so that is addressed in various cryptographic-based papers [Clev, Blum]. Some limitations relate to the capability of cheating parties, and some of them relate to the power of noise over the communication channels [Damgard]. The information-theoretical limits of OT between two parties over a shared channel between three nodes are first addressed in [Hadi-limit] from the same authors. In this paper, we have a more general contribution. As a motivation, we would like to draw attention to the following statement by Kurt Gödel [Godel]:
@@ 4. Information Theoretic Definitions
Let $P_{UVW}$ be a probability distribution over $\mathcal{U}\times\mathcal{V}\times\mathcal{W}$ For any $\epsilon>0$ and $\epsilon^{\prime}>0$ it holds that [Holenstein]:
@@
Also, $H^{\epsilon}_{\infty}(U,V|W)$ can be bounded from below and above as [Holenstein]:
@@
[Holenstein, Corollary 2.12] Let $P_{X^{n}Y^{n}}$ be independent and identically distributed (i.i.d.) according to $P_{XY}$ over the alphabet $\mathcal{X}^{n}\times\mathcal{Y}^{n}$. For any $\epsilon>0$, we have
@@
A function $h:\mathcal{R}\times\mathcal{X}\to\{0,1\}^{n}$ is a two-universal hash function [Carter] if, for any $x_{0}\neq x_{1}\in\mathcal{X}$ and for $R$ uniformly distributed over $\mathcal{R}$, it holds that
@@
[Winter1, Jurg] (Distributed leftover hash lemma) Let $\epsilon>0$, $\epsilon^{\prime}\geq 0$, and let $g_{i}:\mathcal{T}_{i}\times\mathcal{X}_{i}\to\{0,1\}^{n_{i}}$ for $1\leq i\leq m$ be two-universal hash functions. Assume random variables $X_{i}$ over $\mathcal{X}_{i}$, $1\leq i\leq m$, where for any subset $\mathcal{S}\subseteq\{1,2,\dots,m\}$, and $\mathbf{X}|_{\mathcal{S}}=X_{\mathcal{S}(1)},X_{\mathcal{S}(2)},\dots,X_{\mathcal{S}(|\mathcal{S}|)}$, we have
@@
We also briefly note that following from [Cover, Th. 17.3.3], it directly follows from the previously defined $X_{i}$, $T_{i}$, $g_{i}(T_{i},X_{i})$, and $Z$ that
@@
[Bennet, Corollary 4] Let $p_{X,Y}$ be any probability distribution, where $X\in\mathcal{X}$, $Y\in\mathcal{Y}$, and $y$ is a specific realization of $Y$. Assume that $H_{2}(X|Y=y)\geq c$ for some constant $c$. Let $\mathcal{K}$ be a universal class of functions mapping $X$ to $\{0,1\}^{l}$, and let $\kappa$ be sampled uniformly from $\mathcal{K}$. Then:
@@
[Rudolf1, Lemma 3] Let $X$, $Y$, and $Z$ be random variables defined on the finite sets $\mathcal{X}$, $\mathcal{Y}$, and $\mathcal{Z}$, respectively. For any $z_{1},z_{2}\in\mathcal{Z}$ with $p\triangleq\Pr[Z=z_{1}]>0$ and $q\triangleq\Pr[Z=z_{2}]>0$, the following inequality holds:
@@ III-A1 Yao’s Millionaires’ Problem
One of the foundational problems that introduced the concept of secure two-party computation is Yao’s Millionaires’ Problem. In this thought experiment, proposed by Andrew Yao in 1982, two millionaires wish to determine who is richer without revealing their actual wealth to each other. Formally, each party holds a private input—Alice has $x$ and Bob has yy—and they aim to compute the function $f(x,y)=(x>y)$ without disclosing any other information about $x$ or $y$. To solve this problem, Yao introduced the garbled circuit technique, where one party constructs a Boolean circuit representing the function and encrypts (or “garbles”) its components. The other party evaluates the garbled circuit using encrypted inputs obtained via oblivious transfer, ultimately learning only the output. This protocol demonstrated the feasibility of general secure computation between two parties and laid the groundwork for modern cryptographic protocols in this domain [Yao].
@@ III-A2 Blum’s Fair Coin Tossing
Blum’s seminal fair coin tossing (FCT) protocol [Blum] represents one of the earliest cryptographic solutions for enabling two distrustful parties to agree on a random bit in a fair manner. Consider a scenario where Alice and Bob must decide who gets to use their shared home office during a scheduling conflict. Since neither party trusts the other to flip a coin honestly, they employ a cryptographic mechanism to simulate a fair coin toss. Each party independently selects a random value and commits to it using a secure hash function, effectively locking in their choice without revealing it. Once both commitments are exchanged (e.g., over a public channel), they disclose their original values. The XOR or sum of the two values determines the winner: for example, if the result is even, Alice wins; if odd, Bob does. Crucially, the commitment phase prevents either party from altering their input based on the other’s choice, thereby ensuring fairness. If one party refuses to open their commitment or submits an invalid reveal, they automatically lose the toss, which discourages dishonest behavior. This mechanism exemplifies rational fairness in distributed systems, where the optimal strategy for both participants is to follow the protocol faithfully.
Although this game-theoretic interpretation of fairness aligns naturally with rational behavior, it has received comparatively little attention in the mainstream literature on secure MPC [Yao, Yao86, Oded]. Instead, the field has gravitated toward a more rigorous and adversarially robust notion known as unbiasability. Under this definition, the protocol must guarantee that no coalition of malicious participants can skew the outcome of the computation (in particular, a coin toss) regardless of their strategy. Blum’s original coin-tossing scheme does not meet this stronger requirement; while it prevents a player from biasing the outcome in their favor, it does not fully eliminate the possibility of adversarial influence. The concept of unbiasability has been extensively investigated in cryptographic research, where it is well-established that achieving it requires an honest majority among participants [Oded, Rabin2]. Notably, Cleve’s impossibility result [Clev] proves that in the presence of a dishonest minority comprising half or more of the parties, unbiasable coin tossing becomes fundamentally unachievable.
@@ III-A3 Cleve’s Impossibility Result for Fair Coin Tossing
As Cleve’s analysis shows, a malicious Alice has only limited power to influence the outcome, since she does not have access to Bob’s random bits $x_{i}$ during the execution of the protocol. Because of the one-wayness of the trapdoor function $F$, and her inability to invert $F_{K_{i}}(x_{i})$ without the corresponding trapdoor $T_{i}$, she cannot predict or bias the XOR outputs before the trapdoors are revealed.
In contrast, a malicious Bob has greater influence: he may abort the protocol prematurely by withholding the private key $T_{i}$ at some round $i$. In such a case, Alice replaces the unknown values $x_{i},x_{i+1},\ldots,x_{r}$ with random bits, thereby introducing uncertainty into the final majority computation. Nevertheless, Cleve shows that even in the worst case, such selective aborting can alter the majority outcome by only a small amount; specifically, the resulting bias is upper bounded by $O\!\left(\frac{1}{\sqrt{r}}\right)$. He also establishes a lower bound of $\frac{1}{2r}$, demonstrating that a nonzero bias is unavoidable in any finite-round protocol.
Thus, perfect fairness—understood as zero bias—is unattainable for two-party coin-tossing protocols with a finite number of rounds. As the number of rounds $r$ increases, the bias decreases but never completely disappears.
@@ III-B1 In praise of noise
Claude Shannon’s Noisy Channel Coding Theorem [Shannon1] demonstrates that reliable communication is achievable even in the presence of noise, provided the information transmission rate remains below a critical limit known as channel capacity. This foundational result reveals that noise, while disruptive, is not entirely detrimental; rather, it can be mitigated through appropriate encoding schemes and the strategic use of redundancy. Shannon’s insight challenges the notion that randomness and disorder are purely negative phenomena, illustrating that they can be harnessed within structured frameworks to preserve the integrity of information.
@@
In a broader conceptual sense, entropy serves as a metaphor for disorder or chaos in the universe, with far-reaching implications for how we understand complexity and the emergence of order. Information theory’s treatment of noise thus intersects with deeper philosophical considerations about the relationship between chaos and structure. Michel Serres, in particular, explores this theme through the metaphor of the parasite [Serres], suggesting that noise is not merely disruptive but can serve as a generative force. According to Serres, noise introduces feedback and transformation within communicative systems, acting as a catalyst for innovation and the emergence of new meanings. Importantly, one of the most profound insights drawn from noise is its role in enabling the possibility of secure communication. In the sense of OT, now we know why OT can not be obtained from scratch and the existence of noisy resources is a crucial condition for OT, as well as other cryptographic primitives. When no noisy resource is available, the system lacks the necessary randomness to obfuscate the client’s actions from the servers. This absence of noise leads to a situation of complete determinism, where all parties can eventually deduce the hidden information, violating the principles of OT.